OpenStudy (zmudz):

Find the value of \(\theta\), \(0^\circ \le \theta \le 90^\circ\), such that \(\cos 5^\circ = \sin 25^\circ + \sin \theta\). (Give your answer in degrees, not radians.)

OpenStudy (zmudz):

Help much appreciated!

ganeshie8 (ganeshie8):

Hint : sin(25) = sin(90-65) = cos(65)

ganeshie8 (ganeshie8):

Next, to isolate \(\theta\), use below identity CosC - CosD = 2Sin[(C+D)/2]Sin[(D-C)/2]